# Plug-and-Play Transformer Modules for Test-Time Adaptation

### 10 Jan 2024 v2

@@ Abstract
Parameter-efficient tuning (PET) methods such as LoRA, Adapter, and Visual Prompt Tuning (VPT) have found success in enabling adaptation to new domains by tuning small modules within a transformer model. However, the number of domains encountered during test time can be very large, and the data is usually unlabeled. Thus, adaptation to new domains is challenging; it is also impractical to generate customized tuned modules for each such domain. Toward addressing these challenges, this work introduces PLUTO: a Plug-and-pLay modUlar Test-time domain adaptatiOn strategy. We pre-train a large set of modules, each specialized for different source domains, effectively creating a “module store”. Given a target domain with few-shot unlabeled data, we introduce an unsupervised test-time adaptation (TTA) method to (1) select a sparse subset of relevant modules from this store and (2) create a weighted combination of selected modules without tuning their weights. This plug-and-play nature enables us to harness multiple most-relevant source domains in a single inference call. Comprehensive evaluations demonstrate that PLUTO uniformly outperforms alternative TTA methods and that selecting $\leq$5 modules suffice to extract most of the benefit. At a high level, our method equips pre-trained transformers with the capability to dynamically adapt to new domains, motivating a new paradigm for efficient and scalable domain adaptation.
@@ 1 Introduction
Pretrained transformers [7, 4, 31, 38, 39] have achieved significant recent success, however, these models are often quite large, with billions of parameters [4, 39]. This has posed challenges for their practical use, especially in edge devices, due to the computational and memory challenges of fine-tuning for downstream tasks. To address this, parameter-efficient tuning (PET) methods have been proposed including prefix/prompt-tuning [28, 27, 30], adapter [18], and LoRA [20]. In PET, the pretrained model remains frozen and only the domain-specific (extra) parameters designated for the target domain are updated based on the training data. PET has found notable success as it significantly reduces memory usage and facilitates modular adaptation of the model while often being competitive with fine-tuning.
While PET methods have demonstrated notable benefits across various domains and tasks [30, 51], their effectiveness in more challenging scenarios involving few-shot and unlabeled data or domain shifts remain as a major bottleneck [14, 39]. Specifically, we ask: How to leverage PET methods to adapt to a new domain at test-time [52] where we only observe few-shot unlabeled data from this new domain? Here, test-time adaptation (TTA) refers to adapting to new domains during testing, in an online manner, without the need for additional training data. This makes this setting more challenging than standard unsupervised domain adaptation (UDA) [47, 48]. TTA typically requires iteratively updating the source model parameters using an unsupervised objective that incorporates the new test sample from the target distribution. Existing test-time adaptation (TTA) techniques mostly depend on a single source model for adaptation [52, 53]. However, considering the dynamic nature of the test distribution, it should prove advantageous to adapt a set of source models collectively during test time [55, 46].
Table 1: Our work compared to the prior art fulfills key criteria for a comprehensive adaptation framework. To efficiently realize multiple sources, we employ PET modules within a single ViT.
Setting Source Free Adaptation On the fly Dynamic Target Multi Source
@@
Figure 1: The overview of PLUTO. At test time, PLUTO efficiently combines the sources using appropriate weights determined by the current test distribution. Furthermore, we selectively update the LayerNorm (LN) parameters of the model that demonstrates the highest correlation with the test distribution.
@@
In this work, rather than employing multiple distinct source models, we propose to plug source module(s) within a single vision transformer. We naturally can’t expect a single module to work for all target domains. Also, since target domains at test time are few-shot and there can be a large number of domains, it is impractical to train customized PET modules for each target from scratch. Instead, the ideal algorithm should identify the new domain and suitably adapt pre-existing PET modules to this domain. The main contribution of this work is providing a plug-and-play algorithm to address this challenge. We call our method PLUTO: Plug-and-pLay modUlar Test-time adaptatiOn.
@@
Concretely, we consider a multi source domain adaptation setting and propose pretraining a PET module for a collection of source domains. These pretrained modules can then be utilized at test-time with minimal supervision and sample size. Importantly, our algorithm PLUTO benefits from having access to a module store containing a diverse set of source domains. PLUTO can retrieve the most relevant sources from this module store (70+ modules for ImageNet-C) and accurately blend them in a target-aware fashion. This procedure has two important benefits: (1) Sample and compute efficiency: Selecting a small subset of pretrained modules and weighting them is a much lower dimensional search space compared to tuning a module from scratch. (2) Inference efficiency: Unlike alternative multi-source TTA methods such as TENT-Ensemble [52], PLUTO is able to utilize multiple source domains with a single inference call.
@@
Main Contributions. Overall, PLUTO leads to an effective strategy for harnessing not only a pretrained Vision Transformer (ViT) but also pretrained PET modules within this ViT to enhance few-shot unsupervised TTA performance. The main technical contributions of our work are:
@@
  • PET-based TTA. To the best of our knowledge, this is the first work to employ PET for the test-time adaptation problem. We not only pretrain a large number of modules, we present innovative algorithms that can effectively select and blend an optimal subset of these modules to facilitate few-shot unsupervised adaptation to a new domain.
@@
  • Algorithmic innovations. PLUTO facilitates multi-source TTA through weighted logit outputs of multiple modules and fine-tuning of LayerNorms (LN). Our method crucially captures the relationships between the logit outputs of multiple source modules and input instances in a single inference call and enables PLUTO to adaptively decide the benefit of each source domain. We also introduce sharpness-aware minimization (SAM) [10] for TTA and demonstrate how it boosts the TTA performance for PLUTO as well as other TTA/UDA methods.
  • Empirical impact and insights. Comprehensive evaluations demonstrate that PLUTO uniformly outperforms alternative TTA methods and that selecting $\leq$5 modules suffice to extract most of the benefit. We also provide a comparison of zero-shot vs few-shot adaptation which reveals PLUTO also has respectable zero-shot performance.
@@ 2.1 Unsupervised Domain Adaptation
Unsupervised Domain Adaptation (UDA) is widely applied in computer vision tasks, including image classification [48], semantic segmentation [47], object detection [19], and reinforcement learning [40]. These applications address distribution shift challenges by harmonizing source and target domain distributions. Methods utilize techniques like maximum mean discrepancy and adversarial learning [12, 48] for alignment. There is a growing interest in adaptation strategies that rely solely on pretrained source models due to privacy and memory storage concerns regarding source data.
@@ 2.2 Test Time Adaptation (TTA)
UDA demands substantial target data for offline adaptation, while TTA adapts batch by batch. Early work [29] used incoming test-batch stats for adaptation instead of saved training stats. TENT [52] minimizes entropy and updates batch-norm in a pretrained source model with incoming target data. DUA [33] continuously updates batch-norm stats with test batches for target alignment. TTA methods extend to segmentation tasks [49, 44, 21]. However, they often forget source knowledge. CoTTA [53] restores source knowledge stochastically to avoid excessive deviation, and EATA [34] preserves vital model weights with regularization. This paper shows that TTA ensembles mitigate forgetting.
@@ 2.3 Ensemble learning
Ensemble learning, a well-established approach, aims to create a robust and generalizable model [16]. Common ensemble techniques include voting[16], bagging[3], boosting[42, 11], and stacking[54]. These methods have found applications in various tasks, including model debiasing[9, 45], and domain adaptation[25]. Apart from considering global domain-to-domain relationships, [41, 15] introduced the use of similarity measures between examples and domains to choose data from various sources for a particular target domain. These selected data are then merged to train a single classifier. Some closely related works to ours are [25, 37], which also incorporate example-to-domain relationships but utilize an attention mechanism. They train the attention module using limited labeled data from the target domain in a supervised manner. In contrast, our method operates in an unsupervised setting and does not rely on such labeled data.
3 Problem formulation
We are given a large pretrained transformer model $f$ with size $|f|$ and input $x$. Our goal is to develop a domain adaptation technique with "module store", a set of pretrained modules $\{\theta_{j}\}_{j=1}^{N}$ of size $|\theta_{j}|\ll|f|$. Here are the steps building the store $\{\theta_{j}\}_{j=1}^{N}$:
@@ 3 Problem formulation
  • Pretrain on source domains: Given a loss function $\mathcal{L}$ and a fixed set of source domains $\{\mathcal{T}_{j}\}_{j=1}^{N}$, for each $\mathcal{T}\in\{\mathcal{T}_{j}\}_{j=1}^{N}$, instead of fully fine-tuning ($f\rightarrow f_{\mathcal{T}}$) the pretrained model to the domain $\mathcal{T}$ by solving $f_{\mathcal{T}}=\text{argmin}_{f}\mathcal{L}(f;\mathcal{T})$, given the pretrained model (with modules) $f_{\theta}$, we freeze the model and only update the module $\theta=\text{argmin}_{\theta}\mathcal{L}(f_{\theta};\mathcal{T})$.
@@
  • Adapt to a new domain: After obtaining a set of $N$ modules $\{\theta_{j}\}_{j=1}^{N}$ corresponding to $\{\mathcal{T}_{j}\}_{j=1}^{N}$, our objective is to fit a new target domain $\mathcal{T}_{new}$ without fine-tuning a new module but only utilizing the current modules. We aim to find a combination $f$ of the modules $\mathcal{G}(\{\theta_{i}\}_{i=1}^{N})$ that minimizes the loss on the new domain: $\mathcal{G}=\text{argmin}_{\mathcal{G}}\mathcal{L}(f(\mathcal{G}(\{\theta_{j}\}_{j=1}^{N}));\mathcal{T}_{new})$ A very straightfoward version of $\mathcal{G}$ is $\mathcal{G}=\sum_{j=1}^{N}w_{j}\theta_{j}$. An underlying assumption here is that the module output for new target domain $\mathcal{T}_{new}$ can be approximately expressed by the linear combination of current modules outputs.
4 Method
PLUTO’s framework involves four main steps to achieve unlabeled data adaptation:
  • We freeze the weight of the pretrained transformer, and finetune the modules on different source domains in a supervised manner separately.
  • We freeze the weight of the obtained module while keep the pretrained transformer frozen, and train the module selector in a supervised manner.
@@ 4 Method
  • For a test batch at time stamp $t$, we have both the pretrained modules and pretrained models frozen while updating module selector $\mathcal{G}$ with pseudo label entropy minimization. This process helps us learn the module output combination weights and select most related source modules specific to the current batch.
@@
  • At each time stamp $t$, we determine which source module need to update the LN affine parameters by the largest assigned weight from the updated $\mathcal{G}$.
@@
In the second and third steps, we only require a small number of test samples to identify the source modules most relevant to the current test batch. By selecting and using these modules, we achieve results that approximate the effect of using all weighted source module outputs for forward inference. We provided an overview of PLUTO in Alg. 1
4.1 Obtain pretrained modules
@@
Given the correspondence between $N$ source domains $S_{1},\dots,S_{N}$ and their labeled training data, by running parameter-efficient tuning on each domain separately, a set of pretrained modules can be obtained: $\{\theta_{j}\}_{j=1}^{N}$. Given a certain pretrained transformer model $f$, we denote the models equipped with pretrained source modules as $\{f_{\theta_{j}}^{j}\}_{j=1}^{N}$.
For image classification tasks, we assume a shared label space between source and target domains.
@@
1:Input: models with pretrained source modules $\{f_{\theta_{j}}^{j}\}_{j=1}^{N}$, number of modules to be selected $M$ ($M<N$), streaming sequential unlabeled test data $\{x_{i}^{(1)}\}_{i=1}^{B}\rightarrow\{x_{i}^{(2)}\}_{i=1}^{B}\rightarrow\dots\{x_{i}^{(t)}\}_{i=1}^{B}\rightarrow\dots$, Attention module $\mathcal{G}$
2:Output: $M$ scaled weights, $M$ adapted modules
@@
16:     Find indices of top $M$ in $\{w^{*}(t)^{j}\}_{j=1}^{N}$
@@
17:     LN update of source model with largest $\overline{w^{*}(x_{i})^{j}}$ by minimizing Eqn. 4 with Eqn. 8
18:end while
4.2 Initialization and training of module selector

### 08 Feb 2024 v3

@@ Abstract
Parameter-efficient tuning (PET) methods such as LoRA, Adapter, and Visual Prompt Tuning (VPT) have found success in enabling adaptation to new domains by tuning small modules within a transformer model. However, the number of domains encountered during test time can be very large, and the data is usually unlabeled. Thus, adaptation to new domains is challenging; it is also impractical to generate customized tuned modules for each such domain. Toward addressing these challenges, this work introduces PLUTO: a Plug-and-pLay modUlar Test-time adaptatiOn strategy. We pre-train a large set of modules, each specialized for different source domains, effectively creating a “module store”. Given a novel target domain with few-shot unlabeled data, we introduce an unsupervised test-time adaptation (TTA) method to (1) select a sparse subset of relevant modules from this store and (2) create a weighted combination of selected modules without tuning their weights. This plug-and-play nature enables us to harness multiple most-relevant source domains in a single inference call. Comprehensive evaluations demonstrate that PLUTO uniformly outperforms alternative TTA methods across several benchmarks and that selecting $\leq$5 modules suffices to extract most of its benefits. At a high level, our method equips pre-trained transformers with the capability to dynamically adapt to new domains, spearheading a new framework for efficient and scalable domain adaptation.
Machine Learning, ICML
11footnotetext: University of California, Riverside. Email: {cxian008@, sahme047@, bguler@ece., krish@cs., amitrc@ece.}ucr.edu22footnotetext: DEVCOM Army Research Laboratory. Email: ananthram.swami.civ@army.mil33footnotetext: University of Michigan. Email: oymak@umich.edu
@@ 1 Introduction
Table 1: Our work compared to the prior art fulfills key criteria for a comprehensive adaptation framework. We adapt with a sample-specific setup and can support few/zero-shot application.
Setting Source Free Adaptation On the fly Dynamic Target Few/Zero- Shot
@@
Pretrained transformers (Devlin et al., 2018; Brown et al., 2020; Liu et al., 2019; Radford et al., 2019; Raffel et al., 2020) have achieved significant recent success, however, these models are often quite large, with billions of parameters (Brown et al., 2020; Raffel et al., 2020). This has posed challenges for their practical use, especially in edge devices, due to the computational and memory challenges of fine-tuning for downstream tasks. To address this, parameter-efficient tuning (PET) methods have been proposed including prefix/prompt-tuning (Li & Liang, 2021; Lester et al., 2021; Liu et al., 2023), adapter (Houlsby et al., 2019), and LoRA (Hu et al., 2021a). In PET, the pretrained model remains frozen and only the domain-specific (extra) parameters designated for the target domain are updated based on the training data. PET has found notable success as it significantly reduces memory usage and facilitates modular adaptation of the model while often being competitive with fine-tuning.
@@
Figure 1: The overview of PLUTO. At test time, PLUTO efficiently combines the sources using appropriate weights determined by the current test distribution. Furthermore, we selectively update the LayerNorm (LN) parameters of the model that demonstrates the highest correlation with the test distribution. The numbers in the figure are examples provided for illustrative purposes.
While PET methods have demonstrated notable benefits across various domains and tasks (Liu et al., 2023; Vu et al., 2021), their effectiveness in more challenging scenarios involving few-shot and unlabeled data or domain shifts remains a major bottleneck (Gu et al., 2021; Raffel et al., 2020). Specifically, we ask: How to leverage PET methods to adapt to a new domain at test-time (Wang et al., 2020) where we only observe few-shot unlabeled data from this new domain? Here, test-time adaptation (TTA) refers to adapting to new domains during testing, in an online manner, without the need for additional training data. This makes this setting more challenging than standard unsupervised domain adaptation (UDA) (Tsai et al., 2018; Tzeng et al., 2017). TTA typically requires iteratively updating the source model parameters using an unsupervised objective that incorporates the new test sample from the target distribution. Existing test-time adaptation (TTA) techniques mostly depend on a single source model for adaptation (Wang et al., 2020, 2022). However, considering the dynamic nature of the test distribution, it should prove advantageous to adapt a set of source models collectively during test time (Ye et al., 2023; Sun et al., 2015).
@@
In this work, rather than employing multiple distinct source models, we propose to plug source module(s) within a single vision transformer. We naturally cannot expect a single module to work for all target domains. Furthermore, since target domains at test time are few-shot and there can be a large number of domains, it is impractical to train customized PET modules for each target from scratch. Instead, the ideal algorithm should identify the new domain and suitably adapt pre-existing PET modules to this domain. The main contribution of this work is providing a plug-and-play algorithm to address this challenge. We call our method PLUTO: Plug-and-pLay modUlar Test-time adaptatiOn.
@@
Concretely, we consider a multi-source domain adaptation setting and propose pretraining PET modules for a collection of source domains. These pretrained modules can then be utilized at test-time with minimal supervision and sample size. Importantly, our algorithm PLUTO benefits from having access to a module store containing a diverse set of source domains. PLUTO can retrieve the most relevant sources from this module store (70+ modules for ImageNet-C benchmark) and accurately blend them in a target-aware fashion. PLUTO brings three important benefits: (1) Sample efficiency: PLUTO enables few-shot unsupervised adaptation in two ways: First, PLUTO only selects and weights a small subset of pretrained modules, rather than tuning a module from scratch, resulting in an efficient search space. Secondly, we utilize an attention-based module selector to identify the most relevant modules to the target domain. This also allows PLUTO to exhibit strong zero-shot adaptation performance. (2) Anti-forgetting during TTA: Our plug-and-play use of PET modules and selective update of LayerNorm parameters effectively overcomes the forgetting which is a prominent challenge for TTA methods (see Section. 4.5 for details). (3) Parameter efficiency: The adaptation happens through parameter-efficient modules. In a distributed/federated setting where central server operates a module store, the clients can efficiently request new modules.
@@
Main Contributions. Overall, our method PLUTO is an effective strategy for harnessing not only a pretrained Vision Transformer (ViT) but also pretrained PET modules within this ViT to enhance few-shot unsupervised TTA performance. Our main technical contributions are:
@@
  • Multi-source PET-based TTA. To the best of our knowledge, this is the first work to employ multiple PET modules for the test-time adaptation problem (hence the phrase plug-and-play). We not only pretrain a large number of modules, we also present innovative algorithms that can effectively select and blend an optimal subset of these modules to facilitate few-shot unsupervised adaptation to a new domain. Importantly, PLUTO achieves a performance on par with using all modules simultaneously and often outperforms the single best-performing source in hindsight.
@@
  • Algorithmic innovations. PLUTO facilitates multi-source TTA by weighting the output logits of multiple modules and fine-tuning LayerNorms (LN). Our method crucially captures the relationships between the logit outputs of multiple source modules and input instances with very limited samples and enables PLUTO to decide the contribution of each source domain adaptively on the fly. To ensure the generalizability of the few-shot weight assignment of the source modules, we integrate sharpness-aware minimization and PET methods (Foret et al., 2020) for TTA.
  • Empirical impact and insights. Comprehensive evaluations on Digits, Office-Home, CIFAR-10C, and ImageNet-C demonstrate that PLUTO uniformly outperforms alternative TTA methods. The performance improvement is much larger in the few-shot setting. For instance, in Office-Home evaluations in Table 5, PLUTO demonstrate +24% absolute accuracy improvement in zero-shot and 16-shot settings over the best alternative. To do so, PLUTO typically selects only $\leq$5 modules to extract most of the benefits.
@@ 2.1 Unsupervised Domain Adaptation
Unsupervised Domain Adaptation (UDA) is widely used in various machine learning tasks, including image classification (Tzeng et al., 2017), semantic segmentation (Tsai et al., 2018), object detection (Hsu et al., 2020), and reinforcement learning (Raychaudhuri et al., 2021). These applications address distribution shift challenges by harmonizing source and target domain distributions. Methods utilize techniques like maximum mean discrepancy and adversarial learning (Ganin et al., 2016; Tzeng et al., 2017) for alignment. There is a growing interest in adaptation strategies that rely solely on pretrained source models due to privacy and memory storage concerns regarding source data.
@@ 2.2 Test Time Adaptation (TTA)
UDA requires extensive target domain data for offline adaptation, whereas TTA employs a continual, batch-by-batch approach (Valanarasu et al., 2022; Shin et al., 2022; Hu et al., 2021b). Initial studies (Li et al., 2016) adapted using the current test-batch statistics rather than those from training. TENT (Wang et al., 2020) aims to reduce entropy and updates batch-normalization parameters in a pre-trained source model using new target data. DUA (Mirza et al., 2022) persistently updates batch-norm statistics with incoming test batches for better target domain alignment. A common issue with these approaches is the tendency to forget source knowledge. Methods like CoTTA (Wang et al., 2022) apply stochastic source model restoration to avoid drifting of source knowledge, and EATA (Niu et al., 2022) uses regularization to preserve important weights, thus reducing forgetting. In addition, CLIP (Radford et al., 2021)-based methods (Feng et al., 2023; Shu et al., 2022) adapt to new domains by doing data augmentation and prompt generation. These single-source TTA methods require substantial computational resources to balance the trade-off between adaptation performance and mitigating forgetting. In contrast, by assessing the relevance between multiple source models and the target domain and only updating the most related source models, our approach minimizes forgetting with minimal computational expense while maximizing the utilization of knowledge from all source domains.
@@ 2.3 Ensemble learning
Ensemble learning, a well-established approach, aims to create a robust and generalizable model (Hansen & Salamon, 1990). Common ensemble techniques include voting (Hansen & Salamon, 1990), bagging (Breiman, 1996), boosting (Schapire, 1990; Freund & Schapire, 1997), and stacking (Wolpert, 1992). These methods have found applications in various tasks, including model debiasing (Elazar & Goldberg, 2018; Stacey et al., 2020), and domain adaptation (Kim et al., 2017). Apart from considering global domain-to-domain relationships, (Ruder & Plank, 2017; Guo et al., 2018) introduced the use of similarity measures between examples and domains to choose data from various sources for a particular target domain. These selected data are then merged to train a single classifier. Some closely related works to ours are (Kim et al., 2017; Peng et al., 2022), which also incorporate example-to-domain relationships but utilize an attention mechanism. They train the attention module using limited labeled data from the target domain in a supervised manner. In contrast, our method operates in an unsupervised setting and does not rely on such labeled data.
3 Problem Formulation and Proposed Method
We assume access to a pretrained transformer model $f$. We will build a set of pretrained modules, referred to as module store, that can be plugged within $f$ on demand. Let us denote these modules by parameters $\{\theta_{j}\}_{j=1}^{N}$ where each $\theta_{j}$ will be trained for a particular source domain. These modules, based on parameter-efficient tuning methods, will contain much fewer parameters than the base model $f$. Let us also denote the transformer $f$ with module $\theta$ as $f_{\theta}$. Before describing PLUTO, we introduce our workflow: We will (1) pretrain source modules $\{\theta_{j}\}_{j=1}^{N}$ using source datasets and (2) optimally blend these modules for a target task.
@@ 3 Problem Formulation and Proposed Method
  • Pretraining Source Modules: Given a loss function $\mathcal{L}$ and a fixed set of source domains $\{S_{j}\}_{j=1}^{N}$, for each $\mathcal{S}\in\{S_{j}\}_{j=1}^{N}$, instead of fully fine-tuning ($f\rightarrow f^{\mathcal{S}}$) the pretrained model to the domain $\mathcal{S}$ by solving $f^{\mathcal{S}}=\text{argmin}_{f}\mathcal{L}(f;\mathcal{S})$, given the pretrained model (with modules) $f_{\theta}$, we freeze the model and only update the module $\theta=\text{argmin}_{\theta}\mathcal{L}(f_{\theta};\mathcal{S})$.
@@
  • Adapt to a new domain: After obtaining a set of $N$ modules $\{\theta_{j}\}_{j=1}^{N}$ corresponding to $\{\mathcal{S}_{j}\}_{j=1}^{N}$, our objective is to adapt to a new target domain $\mathcal{T}_{new}$ without fine-tuning a new module but only utilizing the current pretrained modules. We aim to find a combination $\mathcal{G}$ of the source models $\mathcal{G}(\{f_{\theta_{j}}\}_{j=1}^{N})$ that minimizes the loss on the new domain:
    $\text{min}_{\mathcal{G}}\mathcal{L}(\mathcal{G}(\{f_{\theta_{j}}\}_{j=1}^{N});\mathcal{T}_{new}).$
    Specifically, $\mathcal{G}$ can be a linear combination of the source models’ output. This would apply under the assumption that the target domain distribution can be expressed as the linear combination of source distributions. We’ll provide some theoretical insights of this approximation in Appendix F.
Method: To proceed, we provide an overview of PLUTO’s framework illustrated in Figure 1. This involves the following steps to achieve unlabeled test-time domain adaptation:
@@
  • (Sec. 3.2) Step 1: For a test batch at time $t$, we have both the pretrained modules and the frozen pretrained transformer model. We update the module selector $\mathcal{G}$ with pseudo-label entropy minimization. This process helps us learn the module output combination weights and select the most related source modules specific to the current batch. As we discuss in Sections 3.1 and 3.2, module selector $\mathcal{G}$ relies on an attention mechanism and plays a crucial role in effectively determining the source modules that are most relevant to the target.
@@
  • (Sec. 3.3) Step 2: At each time $t$, we determine which source model is needed to update the LayerNorm (LN) affine parameters by the largest assigned weight from the updated module selector $\mathcal{G}$.
@@
In both steps, we only require a small number of unlabeled test samples to identify the source modules that are most relevant to the current test batch. We can typically select just a few modules ($\leq$5) to achieve performance that approximates the effect of using all weighted source module outputs during forward inference. The pseudocode for PLUTO is provided in Algorithm 1. In what follows, we further elaborate on the design principles of PLUTO.
3.1 Initialization of the Module Selector
@@
Given the correspondence between $N$ source domains $S_{1},\dots,S_{N}$ and their labeled training data, by running parameter-efficient tuning on each domain separately, a set of pretrained modules, or the module store, can be obtained: $\{\theta_{j}\}_{j=1}^{N}$. Given a certain pretrained transformer model $f$, we denote the model equipped with different pretrained source modules as $\{f_{\theta_{j}}^{j}\}_{j=1}^{N}$.
For source domains $\{S_{j}\}_{j=1}^{N}$ the ground-truth (GT) label of the target sample $x$ is known, and the module selector $\mathcal{G}$ can be initialized by minimizing the cross-entropy loss between the weighted label prediction (will be explained in Sec. 3.2) and the GT label $y$. As discussed in (Peng et al., 2022), this initialization enables $\mathcal{G}$ to capture the sample-specific preference of different source models.
@@
1:Input: Source model with pretrained source modules $\{f_{\theta_{j}}^{j}\}_{j=1}^{N}$, attention-based pretrained module selector $\mathcal{G}$, number of modules to be selected $M$ ($M<N$), streaming sequential unlabeled test data $\{x_{i}^{(1)}\}_{i=1}^{B}\rightarrow\{x_{i}^{(2)}\}_{i=1}^{B}\rightarrow\dots\{x_{i}^{(t)}\}_{i=1}^{B}\rightarrow\dots$
2:Output: $M$ scaled weights, $M$ selected modules, Updated LayerNorm (LN) affine parameters
@@
16:     Find and select the top $M$ in $\{\overline{w}(t)^{j}\}_{j=1}^{N}$
17:     Rescale selected weights to sum up to 1.
@@
18:     LN update of source model corresponding to the largest in $\{\overline{w}(t)^{j}\}_{j=1}^{N}$ by minimizing Eqn. 4 with Eqn. 8
19:end while
3.2 Module weighting at test time